OpenStudy (anonymous):

the equation of the line is y= the order pairs are (0,2) ( 1,7)

OpenStudy (anonymous):

Let's get the slope : \[\text{slope}(f,x,x')= \frac{f(x) - f(x')}{x - x'} \] So here : \[\frac{f(x) - f(x')}{x - x'} = \frac{7 - 2}{1 - 0} = \frac{5}{1} =5 \] Our equation is now something like this : y = 5 x+ b You now have to solve for b with one set of value.